OpenStudy (anonymous):

Write an equation to represent the problem. The area of a triangle is equal to one half times the base times the height. Use A for area, b for base and h for height. A. A = 1/2 b + h B. A = 1/2 (b + h) C. A = (1/2) bh D. A = 1/2 + b + h

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

C ?

jimthompson5910 (jim_thompson5910):

The area of a triangle is equal to one half times the base times the height Area = one half times base times height A = 1/2 times b times h A = (1/2)bh